Allison:

-17.9 - 4.2 Write this in addition form~

Allison:

@Vocaloid

Vocaloid:

I'll be honest, I've never learned the term "addition form" specifically but my best guess would be to just to add the inverse of 4.2 to -17.9

Vocaloid:

so A - B = A + (-B) -17.9 - 4.2 = -17.9 + (-4.2) maybe?

Allison:

Omg. I meant "rewrite the expression using addition".

Vocaloid:

^ yeah I guess it would just be that, then

Allison:

IEY4frWxRviVxlg0sbDJlA.png

Vocaloid:

yeah, that's it

Allison:

You up for answer checking? If you're free~

Vocaloid:

sure

Allison:

Kdf05kAMRFetyTXE240Sxg.png

Vocaloid:

^ yeah that's right

Allison:

_z0aqqHyQLGW7JlXM_gNwA.png

Vocaloid:

yeah that's right

Vocaloid:

apparently if you suspend someone mid-typing it gets stuck like that

Vocaloid:

TIL

Allison:

Oh wow. Who was he? An alt of someone?

Allison:

@wussylover verify your email and talk to people in chat in the bottom right corner and they'll help you.

Allison:

ThRPUsmfTxGu49QJbXKwBw.png

Vocaloid:

yeah, 6.24 is right

Allison:

ufPDcJPER2WTxbqfMIz9qw.png

Vocaloid:

yup, 6.2 is good

Vocaloid:

mrw people think I'm stupid and can't check IP's

Allison:

Olive, did you see my grades? c:

Allison:

Who is it xD

Vocaloid:

yeah I saw, very nice work

Allison:

Thanks for checking my answers ^-^

Allison:

I gotta solve like a few more and in 5 minutes I'll check with you again okay?